An aqueous solution has a hydroxide ion concentration of 1.0×10-8 M   (10-8 M).

(1) What is the hydronium ion concentration in this solution?  M

(2) What is the pH of this solution? 

(3) What is the pOH of this solution?
Answer #1

Sollt. $_Givene Aqueous solution has Onions = 10-8M nt ion concentration = 10-14 Tor = 100M -log Cat pH of solution = 6 of th

answered by: ANURANJAN SARSAM